OpenStudy (cayla):

Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences. The equation is f(x)=175-6x/4 How can I do this?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

it is in function notation already \[f(x)=\frac{175-6x}{4}\] if i am reading it correctly

OpenStudy (cayla):

@misty1212 yes I believe so

OpenStudy (misty1212):

you could rewrite if if you like as \[f(x)=\frac{175}{4}-\frac{3}{2}x\] but is is the same either way

OpenStudy (misty1212):

that way, it is easier to see that the graph is a line with slope \(-\frac{3}{2}\) and \(y\)-intercept \(\frac{175}{4}\)

OpenStudy (cayla):

@misty1212 okay, so two points on my graph would be 0,175 on the y intercept and 150,-50 on the x intercept. How do I explain what the graph of the function represents?

OpenStudy (misty1212):

it would not be \((0,175) \) but rather \((0,\frac{175}{4})\) can you post a screen shot?

OpenStudy (misty1212):

it just seems kind of weird

OpenStudy (cayla):

@misty1212 yeah one sec ill post it

OpenStudy (misty1212):

kk

OpenStudy (cayla):

@misty1212 Here it is I put in the equation 175-6x/4

OpenStudy (misty1212):

oh so it was \[f(x)=175-\frac{6x}{4}\]?

OpenStudy (cayla):

@misty1212 yes

OpenStudy (misty1212):

what i was asking for what the original question, that would help

OpenStudy (misty1212):

ok then that is a line with y intercept \((0,175)\) and slope \(-\frac{3}{2}\)

OpenStudy (seratul):

I think you should make sure what the original equation was so there is no confusion at the end. Parenthesis can make a huge difference in these problems.

OpenStudy (misty1212):

good idea that is why a screen shot would really help

OpenStudy (cayla):

@misty1212 ok sorry for the confusion. I believe I knw what the graph looks like but what does it represent? that is what I don't get and I don't know how to explain that

OpenStudy (misty1212):

i don't know either the graph of the function is the graph of \(y=f(x)\), the points on the graph are all \((x,f(x))\) but i am not sure what the question is actually asking, nor what the function actually is

OpenStudy (the_fizicx99):

Screen shot the question itself, mate

OpenStudy (cayla):

@misty1212 okay, should I consider thinking that the fuction notation is f(x)=175-6x/4 and the graph of the function represents the function? and sure ill add a screenshot

OpenStudy (cayla):

Here is the original question

OpenStudy (the_fizicx99):

'Write the equation in function notation.' I'm guessing you are given one or you are meant to write your own?

OpenStudy (cayla):

The original equation is 6x+4y=700 which turns into slope intercept form which is y=175-6x/4.

OpenStudy (the_fizicx99):

Wow. Lost everything... \( \sf f(x) = y = 175 - \dfrac{6}{4}x \rightarrow 175 - \dfrac{3}{2}x \) Slope is negative, essentially meaning the rate at which the linear function progresses will downscale into larger negative "tones." To find the intercepts, set Y equal to zero and then X equal to zero: 0 = 175 - 3/2x -175/(-3/2) = 116.667 Y-intercept is basic, y = 175 - 3/2(0)